The complex FeSCN^+2 has a molar absorptivity of
Answer #1

Molar absorptivity, \varepsilon = 7.0x103 L/(mol.cm)

Percent transmmittance = 65.9 %

Path length = 1 cm

We have:

Absorbance, A = 2 - log (%T)

= 2 - log (65.9)

= 2 - 1.818

= 0.182

From Beer's law:

A = c*\varepsilon*l

0.182 = c * 7.0x103 * 1

c = 0.026x10-3

c = 2.6x10-5 Mol/L

Hence, concentration = 2.6x10-5 Mol/L